OpenStudy (anonymous):

I really need help with these. they are the ones I couldn't do myself im in a hurry please all the help is much appreciated!!! What is the equation of the line that passes through the points (1,2/3) and (-1,-1) A. Y=-6X+5 B. 5/6X- 1/6 C.Y=2/3 X+1/5 D.Y=6/5X+1/5

OpenStudy (anonymous):

What is the equation of the line that has a slope of -1/3 and a y-intercept of 5/2 A.y=1/3x+ 5/2 B.y=-1/3 x + 5/2 C.y= 5/2 x-1/3 D. y=5/2 x +1/3

OpenStudy (anonymous):

Using point-slope form, write the equation of the line that passes through the point (-1/2,1/2) and has a slope of -1 A. y-1/2=-x-1/2 B.y-1/2=-x+1/2 C.y+1/2=-x-1/2 D.y+1/2=-x+1/2 I think its b but im not forsure I would like some help

OpenStudy (anonymous):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)? A. 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1

OpenStudy (anonymous):

@sloths

OpenStudy (anonymous):

@sloths if you cant figure them out its okay thankyou for looking at them

OpenStudy (anonymous):

haha give me a minute to look them over

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i'm having computer difficulties i'm so sorry this is taking so long

OpenStudy (anonymous):

first question is y+5/6-1/6

OpenStudy (anonymous):

because that's the only one with the correct slope

OpenStudy (math&ing001):

For the first question, plug the 2 points in each equation and see which one checks.

OpenStudy (anonymous):

no your okay if you cant get them its alright thankyou!! I see yeah

OpenStudy (math&ing001):

For the second question, this is how the equation of a line looks like y=mx+b, with m being the slope and b the y-intercept.

OpenStudy (anonymous):

last question is b. can't really explain because i graphed it out haha

OpenStudy (anonymous):

aha its okay thankyou guys so much most people ignore me!

OpenStudy (anonymous):

haha sure whenever i'm online i'll try to help

OpenStudy (math&ing001):

Two lines are parallel when they have the same slope. To check which one passes through the point (2, –1), plug it in all the equations that have the same slope.

OpenStudy (anonymous):

so the second one would be D? or C? im thinking more D than c

OpenStudy (math&ing001):

It's not c or d. This is the form y=mx+b m : slope. b : y-intercept.

OpenStudy (anonymous):

third one is y-1/2=x+1/2

OpenStudy (anonymous):

you just plug the numbers into the point slope formula which is y-y1=m(x-x1)

OpenStudy (math&ing001):

I have to go now. I'll be back later.

OpenStudy (anonymous):

then its b? or am I wrong again

OpenStudy (anonymous):

second one is b

OpenStudy (anonymous):

got you! oh my gosh thank you so much for your help seriously!

OpenStudy (anonymous):

haha sure, any time

OpenStudy (math&ing001):

Yeah b is correct.

OpenStudy (math&ing001):

For the third question, are you sure these are the correct choices : A. y-1/2=-x-1/2 B.y-1/2=-x+1/2 C.y+1/2=-x-1/2 D.y+1/2=-x+1/2 Because D and A are the same and they are both correct.

OpenStudy (anonymous):

D and A are different because A is y minus 1/2 and D is y plus 1/2

OpenStudy (anonymous):

but neither of those are correct, it's b

OpenStudy (math&ing001):

@sloths simplify both and you'll see.

OpenStudy (anonymous):

the formula is y-y1=x-x1 y-1/2=x+1/2

OpenStudy (math&ing001):

And for b, if you plug in (-1/2,1/2) you'll get: 1/2-1/2=-(-1/2)+1/2 0=1

OpenStudy (math&ing001):

The one you gave doesn't figure on the choices.

OpenStudy (math&ing001):

@sloths this the formula y - y1 = m(x - x1) m : the slope which in our case =-1 That would give you y - 1/2 = - (x + 1/2)

OpenStudy (anonymous):

which is still b

OpenStudy (math&ing001):

y - 1/2 = - (x + 1/2) => y - 1/2 = -x - 1/2